### v1

# Threshold Filtering Packing for Supervised Fine-Tuning: Training Related Samples within Packs

Abstract
Packing for Supervised Fine-Tuning (SFT) in autoregressive models involves concatenating data points of varying lengths until reaching the designed maximum length to facilitate GPU processing. However, randomly concatenating data points and feeding them into an autoregressive transformer can lead to cross-contamination of sequences due to the significant difference in their subject matter. The mainstream approaches in SFT ensure that each token in the attention calculation phase only focuses on tokens within its own short sequence, without providing additional learning signals for the preceding context. To address these challenges, we introduce Threshold Filtering Packing (TFP), a method that selects samples with related context while maintaining sufficient diversity within the same pack. Our experiments show that TFP offers a simple-to-implement and scalable approach that significantly enhances SFT performance, with observed improvements of up to 7% on GSM8K, 4% on HumanEval, and 15% on the adult-census-income dataset.
Threshold Filtering Packing for Supervised Fine-Tuning: Training Related Samples within Packs
Jiancheng Dong1, Lei Jiang1, Wei Jin2, Lu Cheng1 1University of Illinois Chicago, 2Emory University dongjiancheng77@gmail.com, ljian43@uic.edu, wei.jin@emory.edu, lucheng@uic.edu
1 Introduction
Figure 1: Overview of TFP. Different from vanilla FT(a), which uses “[PAD]” tokens up to the maximum length, and random packing(b), which places randomly shuffled samples in the same pack and may lead to cross-contamination, TFP(d) places related samples in the same pack while applying a threshold on TSP(c) to ensure that these samples are sufficiently distinct. This approach allows models to learn across diverse contexts and prevent overly similar samples from being grouped together.
In Supervised Fine-Tuning (SFT) for large language models (LLMs), sequence lengths can vary substantially, requiring the wrapping of data into tensors using PyTorch to apply matrix operations optimized for CUDA and GPUs Raffel et al. (2023). As illustrated in Figure 1(a), vanilla fine-tuning pad shorter sequences with special tokens "[PAD]" up to the maximum sequence length. While this ensures uniformity, it introduces inefficiencies by including irrelevant padding tokens in the computation, wasting GPU resources and dilutes the model’s learning signal Kundu et al. (2024).
To address this issue, packing sequences has become a common technique in autoregressive transformer models during training and inference to optimize context length and reduce padding Liu et al. (2019); Brown et al. (2020). This method involves randomly selecting and concatenating data of varying lengths until reaching the designed maximum length. Recent studies suggest that packed data, when batched and processed on multi-GPUs, effectively minimize idle time within each batch Bai et al. (2024).
However, randomly concatenating these data samples, which are then fed into an autoregressive transformer (Figure 1(b)), can result in sequence cross-contamination Krell et al. (2022). This occurs when predictions for one sequence are influenced by an unrelated sequence, making accurate predictions challenging, especially when the sequences pertain to different subjects. An example of cross-contamination is when a model is instructed to generate a multiplication table, followed by an instruction to create a list of useful expressions in French. If these two sequences are concatenated without careful separation, the model might incorrectly mix instructions and outputs, such as generating multiplication results interspersed with French phrases like "3 x 2 = Bonjour," leading to inappropriate and confusing outcomes. Moreover, current SFT pipelines Kundu et al. (2024) cause previous samples to provide no signal for predicting the next sample, thereby reducing learning efficiency and negatively impacting the few-shot performance of LLMs.
To address these challenges, in this work, we present Threshold Filtering Packing (TFP), a new packing approach that packs sequences of related yet diverse samples, encouraging context richness and reasoning across sample boundaries. Specifically, we employ a greedy algorithm inspired by the Traveling Salesman Problem (TSP) Applegate et al. (2006) to efficiently map out a path for segmentation into multiple packs. TFP is implemented to further refine these packs by ensuring that overly similar samples are not grouped together. Setting the threshold is crucial in this context as it allows us to strike a balance between similarity and diversity within each pack, preventing homogeneity and ensuring the robustness of the generated packs. As shown in Figure 1(c), each sample is first converted to an embedding and then represented as a node in the graph. As shown in Figure 1(d), TFP employs threshold filtering to ensure each sample is distinct enough from recent ones, preventing the model from merely replicating previous outputs. This method results in packs that provide useful context while avoiding cross-contamination from unrelated texts.
For experiments, we fine-tune various LLMs on standard instruction fine-tuning datasets and conduct bias-related experiments to evaluate the potential effect of packing on bias amplification. Additionally, we assess the impact of TFP on computational efficiency for SFT. Our findings indicate that TFP not only demonstrates superior performance across various models but also substantially lowers computational costs. The bias-related experiments show that TFP offers a flexible operational space, allowing for adjustments in the ratio of sensitive attributes (e.g., race) within packs to effectively manage bias.
2 Related Work
SFT and Alignment Fine-tuning is a prevalent strategy to enhance model performance on downstream tasks, evidenced in domains such as coding Wei et al. (2023); Luo et al. (2023) and arithmetic Xiang Yue (2023). Other work has highlighted the importance of consistency in format Liang et al. (2024), data quality Chung et al. (2022), and mixing tasks from different categories Longpre et al. (2023); Iyer et al. (2023) in SFT. As LLMs evolve, the risk of generating unsafe content increases Su et al. (2024); Wang et al. (2024a). Established methods for LLM alignment include instruction fine-tuning and reinforcement learning with human feedback (RLHF) Ouyang et al. (2022). Instruction fine-tuning, also known as SFT, refines pre-trained models using annotated instructional data, often preceding RLHF to aid initial alignment Touvron et al. (2023). RLHF employs reinforcement learning to adapt models based on feedback on generated responses. Although RLHF has been pivotal for developing systems like ChatGPT OpenAI (2021), isolated instruction fine-tuning can yield comparable outcomes Sun et al. (2023) with much less computational and labor costs.
Packing While packing is relatively less researched, it is a technique extensively used in frameworks like Hugging Face’s SFT Trainer  [^1] https://huggingface.co/docs/trl/sft_trainer to expedite inference and training. To prevent cross-contamination during self-attention calculation, existing packing approaches involve concatenating sequences into a single tensor and using masking to disregard elements from other sequences during computation Kundu et al. (2024). This method, including variations like LongAlign Bai et al. (2024) and Prepacking Zhao et al. (2024), enhances training efficiency and minimizes the cross-contamination impact on model performance. However, it necessitates calculating a distinct attention mask for each batch, complicating implementation and increasing memory consumption for masks, which can hinder the effectiveness of flash attention.
Our method differs from previous packing approaches that require masking to prevent cross-contamination. Instead, TFP creates packs of data points that provide useful context without the need for additional masking, simplifying implementation and reducing memory overhead.
3 Threshold Filtering Packing
The standard practice in packing is to form a pack by concatenating random samples until reaching the maximum context length Zhao et al. (2024). However, randomly concatenated packs do not provide additional learning signals and can lead to cross-contamination of sequences, compared to training on individual samples. In contrast, TFP generates more coherent packs by concatenating related and useful samples together, improving SFP performance and computational efficiency.
3.1 Problem Statement
Given a set of samples $D=\{d_{i}\}$, where each sample $d_{i}$ has its instruction converted into an embedding $e_{i}$, our goal is to organize these samples into packs such that each of them consists of related samples that provide semantic context. Formally, we aim to form a set of packs $P_{1}\cdots P_{m}$ where each pack $P_{i}=\{d_{1},\ldots,d_{k_{i}}\}\subset D$ and $\bigcup_{i=1}^{m}P_{i}=D$.
3.2 $k$-NN Packing
An intuitive method for packing is to apply $k$-NN and place each sample along with its retrieved top-$k$ neighbors in the same pack, referred to as $k$-NN packing. This approach maintains sample similarity within each pack but introduces a significant issue: data repetition. Some samples frequently appear as nearest neighbors for multiple others, leading to overlapping packs, i.e., $\exists i\neq j,P_{i}\cap P_{j}\neq\emptyset$. For instance, in the codealpaca2k dataset Chaudhary (2023), the sample "Construct a loop in Python to display all elements in a list" was included in 94 different packs with $k$-NN Packing, greatly reducing the diversity of pack content.
The data repetition problem can contaminate both individual packs and the entire training process. Within a pack, popular samples that are close to many others in the embedding space do not serve as diverse contexts, increasing the risk of cross-contamination. Across the training process, repeated exposure to these popular samples reduces the diversity of the dataset, potentially leading to overfittingShi et al. (2024).
3.3 Threshold Filtering Packing Algorithm
To address these challenges, we propose packing data samples that provide meaningful context while avoiding repeated selection. A basic approach involves using a greedy algorithm to select samples with the smallest Euclidean distance to their corresponding embeddings, ensuring each sample is included only once. This is essentially a greedy algorithm for the TSP Applegate et al. (2006). Intuitively, $k$-NN can select the same data point multiple times when retrieving the nearest neighbors, whereas TSP selects each data point only once.
Further, previous studies Yasunaga et al. (2023); Liu et al. (2024) show that maintaining diversity among the input contexts is crucial. We adopt a greedy algorithm for TSP with conditional adjustments and then segment this path into multiple packs to generate packs composed of diverse and relevant samples. TFP is designed to assemble related samples, with Threshold Filtering specifically addressing the challenge of placing overly similar samples in the same pack. As shown in Figure 1(d), threshold filtering separates overly similar embeddings, such as embedding1 and embedding2, which were initially connected by TSP.
The mathematical formulation of this approach is as follows:
$\text{Minimize}\quad\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\sum_{l=j+1}^{k_{i}}w_{e_{ij},e_{il}}$
subject $\text{subject to}\quad\bigcup_{i=1}^{m}P_{i}=D,$
$P_{i}\cap P_{j}=\emptyset\quad\forall i\neq j,$
$|P_{i}|=k_{i}\quad\forall i,$
$w_{ij}=\sqrt{\sum_{q=1}^{n}(d_{i,q}-d_{j,q})^{2}},$
$w_{e_{ij},e_{il}}>t\quad\text{for all pairs of recent samples}$
Here, $m$ represents the number of packs. $k_{i}$ is the number of elements in the $i$-th pack. $e_{ij}$ refers to the $j$-th element in the $i$-th pack. $w_{ij}$ is the Euclidean distance between elements $e_{i}$ and $e_{j}$. $d_{i,k}$ represents the $k$-th feature of element $e_{i}$. $n$ is the dimensionality of the feature space. $t$ is the distance threshold to ensure diversity within a pack.
This objective function minimizes the sum of the Euclidean distances $w_{e_{ij},e_{il}}$ between all pairs of elements within each pack $P_{i}$. These constraints guarantee that the packs are disjoint, meaning no two packs share any common elements, and ensure that each pack $P_{i}$ contains exactly $k_{i}$ elements. The Euclidean distance between elements $e_{i}$ and $e_{j}$ is calculated based on their feature vectors $d_{i,k}$. Additionally, to ensure diversity within each pack, the constraint $w_{e_{ij},e_{il}}>t$ is applied . This method results in packs that provide useful context while avoiding cross-contamination from unrelated texts.
Algorithm 1 Threshold Filtering Packing
1:  Input: Instruction embeddings array $E$, distance threshold $t$, the distance threshold number $r$
2:  Output: A shortest path satisfying the threshold condition
3:  Initialize $n\leftarrow\text{length of }E$, $visited\leftarrow[\text{False}]\times n$ ,$visited[0]\leftarrow\text{True}$
4:  for $i\in\{1,2,\ldots,n-1\}$ do
5:     $next\leftarrow$ SelectNext($E$, $visited$, $t$, $r$)
6:     // SelectNext($E$, $visited$, $t$, $r$) select the nearest unvisited embedding in $E$ while ensuring the distance from the recent $r$ samples is greater than $t$
7:     $path\leftarrow path\cup[next]$
8:     $visited[next]\leftarrow\text{True}$
9:  end for
10:  return  $path$
As depicted in Algorithm 1, TFP consists of three primary steps: first, generating sentence embeddings $E$ for the instruction parts of the samples, then selecting the nearest unvisited embedding in $E$ while ensuring that the distance from the recent $r$ samples is greater than distance threshold $t$. Subsequently, we use the pack formed by instruction-related samples to train the language model. Since TFP only changes the distribution of data within the pack, it can be seamlessly integrated into existing SFT pipelines for LLMs. As a final step, we traverse the samples along the path and concatenate them to create packs.
4 Experiment
Llama2-7B Llama3-8B Mistral-7B
Method WR HumanEval GSM8K WR HumanEval GSM8K WR HumanEval GSM8K
Vanilla FT 48.2 $\pm$ 0.4 19.5 $\pm$ 0.3 26.2 $\pm$ 0.0 51.2 $\pm$ 0.5 38.4 $\pm$ 0.0 61.8 $\pm$ 0.2 62.9 $\pm$ 0.6 35.4 $\pm$ 0.3 59.7 $\pm$ 0.5
Sorted batching 48.2 $\pm$ 0.5 20.1 $\pm$ 0.3 26.5 $\pm$ 0.0 51.8 $\pm$ 0.6 37.2 $\pm$ 0.4 62.0 $\pm$ 0.6 61.2 $\pm$ 0.3 34.1 $\pm$ 0.5 61.0 $\pm$ 0.1
Random packing 47.1 $\pm$ 0.3 19.5 $\pm$ 0.3 26.1 $\pm$ 0.4 51.7 $\pm$ 0.5 37.8 $\pm$ 0.3 62.1 $\pm$ 0.6 62.4 $\pm$ 0.0 34.8 $\pm$ 0.6 59.1 $\pm$ 0.4
Random packing (mask) 47.6 $\pm$ 0.5 19.5 $\pm$ 0.6 26.1 $\pm$ 0.2 52.4 $\pm$ 0.0 37.8 $\pm$ 0.3 62.5 $\pm$ 0.4 63.5 $\pm$ 0.2 35.4 $\pm$ 0.5 59.2 $\pm$ 0.1
Packing+loss weighting 47.1 $\pm$ 0.6 18.9 $\pm$ 0.4 25.8 $\pm$ 0.0 51.2 $\pm$ 0.3 38.4 $\pm$ 0.0 60.9 $\pm$ 0.3 60.6 $\pm$ 0.4 34.8 $\pm$ 0.5 59.5 $\pm$ 0.0
$k$-NN packing 45.3 $\pm$ 0.0 15.9 $\pm$ 0.6 29.3 $\pm$ 0.3 48.8 $\pm$ 0.4 36.0 $\pm$ 0.0 59.5 $\pm$ 0.5 55.3 $\pm$ 0.4 34.1 $\pm$ 0.3 57.2 $\pm$ 0.2
TFP 51.2 $\pm$ 0.2 22.6 $\pm$ 0.3 33.6 $\pm$ 0.4 54.1 $\pm$ 0.6 42.7 $\pm$ 0.5 66.7 $\pm$ 0.3 63.5 $\pm$ 0.0 38.4 $\pm$ 0.5 64.1 $\pm$ 0.4
Table 1: Comparison of different methods and Different Training Data: Alpaca, CodeAlpaca-2k, and GSM8K, with results represented by WR (Win Rate), HumanEval, and GSM8K, respectively. In the table, we follow the most common few-shot settings: using Win rate judged by PandaLM and a 0-shot setting for HumanEval, while the GSM8K task uses a 4-shot setting.
We evaluate the proposed approach from three aspects: (1) Performance Comparison: We compare TFP with various LLMs fine-tuned on common instruction fine-tuning datasets under both zero-shot and few-shot settings. (2) Bias and Fairness: Inspired by previous research Wang et al. (2024a) that studies the impact of the ratio of different demographic groups in in-context learning (ICL) on LLM fairness, we investigate how adjusting the ratio in each pack during SFT can influence the bias and fairness of LLMs. (3) Efficiency: We study how various SFT methods influence computational efficiency on different GPU setups.
4.1 Experimental Setup
Datasets. We use commonly adopted datasets for instruction fine-tuning, which include tasks related to helpfulness, code-generation capabilities, and mathematical reasoning: (1) Alpaca dataset, generated from the Self-Instruct method Wang et al. (2023b) via the text-davinci-003 model Buruk (2023), covering various tasks such as arithmetic, coding, and question-answering. (2) Code Alpaca datasetChaudhary (2023), which aims to build and share an instruction-following LLaMA model for code generation.(3) GSM8K dataset Cobbe et al. (2021), curated to examine mathematical reasoning capabilities, comprises 8.8k high-quality arithmetic word problems designed at the grade school level.
For the fairness-related experiments, we use the Jigsaw Unintended Bias in Toxicity Classification task cjadams (2019) and Adult dataset Becker and Kohavi (1996). The Jigsaw Unintended Bias in Toxicity Classification task involves performing toxicity classification on comment texts published by the Civil Comments platform. It contains human-annotated demographic information such as race, gender, and religion. The goal is to ensure that models make predictions based on the text toxicity, rather than demographic information included in the text. We use race (Black and non-Black) as the protected attributes. The Adult dataset is a tabular dataset that includes 14 attributes of a person (e.g., age and education level) as input, to predict whether the person’s income exceeds \$50k per year. We evaluate the fairness of fine-tuned models based on the sensitive attribute of sex, specifically comparing “male” and “female”.
Baselines. We compare TFP with the following baselines:
(1) Vanilla fine-tuning: This method appends special padding tokens to shorter prompts to match the maximum length within a batch. Huggingface’s inference framework Wolf et al. (2020) generates corresponding attention masks to ensure the language model disregards the padded tokens during computation to handle prompts of variable lengths.
(2) Sorted batching Bai et al. (2024): This approach sorts inputs by length and samples batches to minimize padding. As a result, each batch consists entirely of either long or short sequences.
(3) Random packing: Random packing involves concatenating data of varying lengths randomly until reaching the maximum length Brown et al. (2020).
(4) Random packing (mask): A variant of random packing that uses masking to prevent cross-contamination between different sequences within the same pack during self-attention calculations.
(5) Packing + loss weighting Bai et al. (2024): A typical packing strategy skews towards longer sequences and those with more target tokens, as packs with fewer sequences or more target tokens disproportionately influence the final loss, especially for datasets designed for long contexts. This method ensures equal loss weighting for each sequence.
(6) $k$-NN packing: In this method, each sample is directly placed together with its retrieved top-k samples in the same pack.
Evaluation Metrics. We follow commonly used protocols Luo et al. (2024); Xiang Yue (2023); Ge et al. (2024) to evaluate SFT in LLMs. Specifically, we use PandaLM Wang et al. (2023a, 2024b) to evaluate the helpfulness of various models. PandaLM provides reproducible and automated comparisons between different LLMs. By providing PandaLM with the same context, it can compare the responses of different LLMs, offer reasons for the decisions, and provide a reference answer. We report the win rate (WR), which is the proportion of instances where the responses are favored over those produced by GPT-3.5 Brown et al. (2020). The code generation skills are enhanced using the Code Alpaca dataset Chaudhary (2023), while evaluation is conducted using the HumanEval dataset Chen et al. (2021). GSM8K dataset Cobbe et al. (2021) uses its own test set. We followed the most common few-shot settings: using Win rate judged by PandaLM and a 0-shot setting for HumanEval, while the GSM8K task uses a 4-shot setting.
We utilize the Llama2-7B Touvron et al. (2023), Llama3-8B AI@Meta (2024), and Mistral-7B Jiang et al. (2023) as the base LM in our experiments. Due to limited computation resources, we employ the QLoRA technique Dettmers et al. (2023) in all fine-tuning experiments. To ensure fair comparison, we maintain consistency in nearly all hyperparameters across all methods. For all results below, we run the experiments five times and report the mean and standard deviations for all compared methods.
4.2 Results
Method 0-shot 4-shot 32-shot
ACC $\uparrow$ $M_{dpd}$downarrow $M_{eod}$downarrow ACC $\uparrow$ $M_{dpd}$$\downarrow$ $M_{eod}$$\downarrow$ ACC $\uparrow$ $M_{dpd}$downarrow $M_{eod}$$\downarrow$
Vanilla FT 0.84 $\pm$ 0.01 0.10 $\pm$ 0.00 0.12 $\pm$ 0.01 0.75 $\pm$ 0.02 0.08 $\pm$ 0.01 0.10 $\pm$ 0.02 0.73 $\pm$ 0.01 0.08 $\pm$ 0.00 0.16 $\pm$ 0.02
Random packing 0.78 $\pm$ 0.02 0.10 $\pm$ 0.01 0.14 $\pm$ 0.02 0.73 $\pm$ 0.02 0.10 $\pm$ 0.01 0.14 $\pm$ 0.02 0.74 $\pm$ 0.02 0.10 $\pm$ 0.01 0.14 $\pm$ 0.02
Random packing (mask) 0.84 $\pm$ 0.00 0.08 $\pm$ 0.01 0.12 $\pm$ 0.02 0.71 $\pm$ 0.02 0.08 $\pm$ 0.01 0.12 $\pm$ 0.02 0.72 $\pm$ 0.01 0.08 $\pm$ 0.01 0.12 $\pm$ 0.02
Balanced ratio 0.71 $\pm$ 0.02 0.04 $\pm$ 0.01 0.08 $\pm$ 0.01 0.57 $\pm$ 0.02 0.06 $\pm$ 0.01 0.12 $\pm$ 0.00 0.64 $\pm$ 0.02 0.03 $\pm$ 0.01 0.06 $\pm$ 0.01
Resampling 0.85 $\pm$ 0.01 0.11 $\pm$ 0.02 0.16 $\pm$ 0.00 0.66 $\pm$ 0.02 0.07 $\pm$ 0.01 0.14 $\pm$ 0.02 0.71 $\pm$ 0.02 0.14 $\pm$ 0.01 0.28 $\pm$ 0.02
TFP 0.87 $\pm$ 0.01 0.06 $\pm$ 0.01 0.10 $\pm$ 0.02 0.77 $\pm$ 0.02 0.10 $\pm$ 0.01 0.24 $\pm$ 0.02 0.81 $\pm$ 0.01 0.01 $\pm$ 0.01 0.02 $\pm$ 0.00
TFP (Balanced) 0.85 $\pm$ 0.01 0.04 $\pm$ 0.01 0.10 $\pm$ 0.01 0.78 $\pm$ 0.01 0.01 $\pm$ 0.00 0.04 $\pm$ 0.01 0.80 $\pm$ 0.01 0.01 $\pm$ 0.01 0.02 $\pm$ 0.01
TFP (Resampling) 0.87 $\pm$ 0.01 0.08 $\pm$ 0.01 0.12 $\pm$ 0.01 0.78 $\pm$ 0.01 0.06 $\pm$ 0.01 0.12 $\pm$ 0.01 0.83 $\pm$ 0.01 0.05 $\pm$ 0.01 0.06 $\pm$ 0.00
Table 2: Accuracy and group fairness metrics on Llama3-8B for the Jigsaw Dataset
Method 0-shot 4-shot 32-shot
ACC $\uparrow$ $M_{dpd}$downarrow $M_{eod}$downarrow ACC $\uparrow$ $M_{dpd}$downarrow $M_{eod}$$\downarrow$ ACC $\uparrow$ $M_{dpd}$$\downarrow$ $M_{eod}$downarrow
Vanilla FT 0.78 $\pm$ 0.02 0.26 $\pm$ 0.03 0.44 $\pm$ 0.04 0.67 $\pm$ 0.02 0.06 $\pm$ 0.01 0.09 $\pm$ 0.02 0.54 $\pm$ 0.03 0.04 $\pm$ 0.01 0.08 $\pm$ 0.02
Random packing 0.76 $\pm$ 0.03 0.25 $\pm$ 0.02 0.42 $\pm$ 0.04 0.50 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.50 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Random packing (mask) 0.78 $\pm$ 0.02 0.25 $\pm$ 0.02 0.44 $\pm$ 0.04 0.65 $\pm$ 0.02 0.06 $\pm$ 0.01 0.08 $\pm$ 0.02 0.56 $\pm$ 0.02 0.06 $\pm$ 0.01 0.08 $\pm$ 0.02
Balanced ratio 0.78 $\pm$ 0.02 0.28 $\pm$ 0.03 0.40 $\pm$ 0.03 0.64 $\pm$ 0.02 0.03 $\pm$ 0.01 0.06 $\pm$ 0.02 0.56 $\pm$ 0.02 0.02 $\pm$ 0.01 0.02 $\pm$ 0.01
Resampling 0.72 $\pm$ 0.03 0.21 $\pm$ 0.02 0.36 $\pm$ 0.03 0.58 $\pm$ 0.03 0.04 $\pm$ 0.01 0.06 $\pm$ 0.02 0.50 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
TFP 0.80 $\pm$ 0.02 0.22 $\pm$ 0.02 0.32 $\pm$ 0.03 0.81 $\pm$ 0.02 0.08 $\pm$ 0.01 0.10 $\pm$ 0.02 0.78 $\pm$ 0.02 0.02 $\pm$ 0.01 0.04 $\pm$ 0.02
TFP (balanced) 0.80 $\pm$ 0.02 0.11 $\pm$ 0.02 0.16 $\pm$ 0.03 0.78 $\pm$ 0.02 0.02 $\pm$ 0.01 0.02 $\pm$ 0.01 0.74 $\pm$ 0.02 0.01 $\pm$ 0.01 0.02 $\pm$ 0.01
TFP (resampling) 0.81 $\pm$ 0.02 0.10 $\pm$ 0.01 0.16 $\pm$ 0.02 0.80 $\pm$ 0.02 0.14 $\pm$ 0.02 0.18 $\pm$ 0.03 0.79 $\pm$ 0.02 0.15 $\pm$ 0.02 0.24 $\pm$ 0.03
Table 3: Accuracy and group fairness metrics on Llama3-8B for the Adult Dataset
Comparisons of Various Packing Methods in Instruction Fine-tuning
Table 1 displays the results of fine-tuning on three downstream datasets: Alpaca Taori et al. (2023), codealpaca-2k Chaudhary (2023), and GSM8K Cobbe et al. (2021). We have the following key observations:
(1) TFP consistently outperforms the best of other baselines across various base LLMs, achieving improvements of up to 7% on GSM8K, 4% on HumanEval, and 3% on Alpaca. This suggests that TFP learns more effectively about answering questions from the related data within the pack. However, naive packing strategies (random packing, random packing (mask)) fail to improve performance and can even degrade it.
(2) Comparing TFP with $k$-NN packing, we find that $k$-NN packing can lead to performance declines, especially in well-trained models like Llama3-8B and Mistral-7B, whereas TFP consistently outperforms. Although Llama2-7B showed some improvement on GSM8K with $k$-NN packing, as inferred from previous work Yuan et al. (2023), these gains are likely due to repeated training effects rather than genuine generalization. In contrast, TFP enhances diversity and relevance within packs, leading to better performance without overfitting, underscoring its robustness.
(3) Sorted batching and loss weighting methods, which are designed for long context, prove less effective on shorter datasets like GSM8K, where the average sequence length of around 500 tokens is far below the model’s maximum length. As highlighted in previous work Bai et al. (2024), sorted batching can introduce bias in data distribution across batches, where entire batches consist of either long or short sequences, potentially disrupting the optimization process during stochastic gradient descent.
Impact of TFP on Fairness
2*L40s 1*3090
Method LLaMA2 7B LLaMA3 8B Mistral 7B LLaMA2 7B LLaMA3 8B Mistral 7B
Vanilla FT 1.73 1.05 1.15 4.68 4.43 5.25
Sorted batching 1.70 1.05 1.08 4.68 4.42 5.22
Random packing 0.37 0.40 0.40 2.53 2.03 2.58
Random packing (mask) 0.40 0.38 0.40 2.57 2.50 2.62
Packing + loss weighting 0.42 0.43 0.45 2.48 2.07 2.55
$k$-NN packing 1.57 1.35 1.42 9.68 9.40 9.87
TFP 0.40 0.37 0.38 2.57 2.02 2.55
Table 4: Training time (hrs) of LLaMA3-8B using different methods on GSM8K.
Our method, which packs similar data together in the same pack, may lead to imbalanced data across different demographic groups, potentially amplifying biases. We explore two approaches: TFP (balanced) and TFP (resampling) to achieve balance while maintaining text relevance. We also use balanced ratio and resampling methods on the data’s default sequence as controls. Particularly, inspired by DecodingTrust Wang et al. (2024a) that shows a balanced ratio across different groups in ICL can help improve LLM fairness, we investigate whether a balanced ratio of different demographic groups within a pack during SFT influences LLM fairness using classification tasks. We report prediction accuracy (ACC) and two common fairness metrics used for classification, equalized odds differences (eod) Hardt et al. (2016) and demographic parity difference (dpd) Zemel et al. (2013). Their detailed definitions can be found in Appendix A.
We use 0-shot and balanced 32-shot ICL settings (i.e., the ratio of different groups in ICL is 1:1) for evaluation following the experimental design of DecodingTrust Wang et al. (2024a), and we also explore the effects of a small number of samples with a balanced 4-shot setting. Experiments in Tables 2 and 3 were conducted on text and tabular datasets respectively, from which we have the following three observations.
(1) TFP (balanced) excels in fairness tasks for both text and tabular data, especially when the number of ICL examples is small. When the number becomes large (e.g., 32), the influence of the balanced ratio in ICL examples on fairness becomes dominant. The original TFP does not excel in fairness in 0-shot settings due to imbalanced data across social groups. TFP (resampling) results in instability due to the repeated sampling of certain data. By providing the model with hard negative examples (i.e. closely positioned samples with differing labels) within a pack, TFP enables more efficient learning from data with similar texts but different labels. Balancing the ratio within packs introduces many samples with similar texts but varying sensitive attributes, which helps mitigate biases in LLMs.
(2) In nearly all settings, TFP and its variants demonstrate superior accuracy, particularly on tabular data (as shown in Table 3). LLMs are known not to excel in prediction tasks for tabular data, especially under ICL settings Fang et al. (2024). When using an excessive number of ICL examples, all baseline approaches tend to predict the same value for all samples, showing degraded prediction performance. TFP, however, presents strong potential in ICL, with improved fairness and competitive accuracy (up to 15% improvement in accuracy).
(3) Conventional packing methods often struggle to balance the trade-off between fairness and accuracy. For example, when the ratio within a pack is adjusted to achieve fairness, accuracy tends to decrease across all shot settings. Additionally, the instability caused by repeated training through direct resampling is more pronounced compared to TFP (resampling), possibly because unrelated sequences can disrupt the model’s judgment.
Figure 2: Training time (hrs) on 2*L40S 48G and 1*3090 under different training methods.
4.3 Impact of Packing on Efficiency
Previous work has proposed two methods for handling long data: sorted batching and packing with loss weighting Bai et al. (2024). It is suggested that these two methods can reduce idle time and speed up the training process across multiple GPUs. The acceleration achieved by these two methods on long data is approximately the same.
In this experiment, we study how various SFT methods influence computational efficiency on different GPU setups. We select the GSM8K dataset due to its widespread use and report the SFT time of each approach in Figure 2. The results under different GPU settings are reported in Table 4.
Figure 3: Performance comparison of different embedding models and different data segments on GSM8K.
Figure 4: Accuracy of TFP on Llama Models of different distance threshold number $r$.
We observe that: (1) When data lengths are much shorter than the maximum token limit of LLMs, packing significantly reduces training costs, particularly in the fine-tuning phase, by optimizing CUDA matrix operations. This improvement is beneficial for both multi-GPU and single-GPU configurations. (2) Our packing method reduces the need for distinct attention masks per batch by allowing a single standard diagonal for the whole batch. It avoid increased mask memory consumption from incorrect implementation.
5 Ablation Studies
5.1 TFP Design
We conducted a series of ablation studies on several critical design choices for TFP using the GSM8K dataset due to its widespread use and high quality. Embedding models play a pivotal role in these ablation studies, as they capture different semantic meanings, which is essential for understanding the operational mechanisms of TFP.
Initially, we evaluated various embedding models, including bert-base-uncased Devlin et al. (2019), RoBERTa Liu et al. (2019), all-MiniLM-L6-v2, E5-mistral-7b-instruct Wang et al. (2022), and the hidden layers of Llama3 AI@Meta (2024). These models fall into two primary categories Liu et al. (2024): the Model-based method, which includes base models like bert-base-uncased, Llama3, and RoBERTa, and the Semantic-based method, which employs advanced sentence embedding models such as E5-mistral-7b-instruct and all-MiniLM-L6-v2.
As presented in Figure 4, the base models generally outperformed the others, with bert-base-uncased emerging as the most effective. This result suggests that base models are better suited for sample aggregation in contextual training, likely because TFP requires models that provide contextual information rather than those focused solely on semantic similarity as in clustering or retrieval tasks.
Further analysis explored the optimal data segments for TFP within the Alpaca format, which typically includes instruction, input, and output components. Given that inputs often lack substantial content, we conducted experiments on the instruction part, output part, and entire data. Our experiments on the GSM8K dataset indicated that using instructions for similarity calculations was the most beneficial. This finding is significant because calculating similarity within the instruction section reduces the risk of cross-contamination. By avoiding the clustering of similar outputs, we can prevent models from simply reproducing previous outputs since SFT only calculates loss on the outputs.
5.2 The distance threshold number $r$
We now examine the influence of the distance threshold number $r$ on the TFP algorithm. The parameter analysis were conducted using Llama3-8B and Llama2-7B to determine how variations in $r$ impact performance within models from the same series. The results, evaluated on the GSM8K dataset, are presented in Figure 4.
Focusing solely on $r$, setting $r=0$ causes a greedy selection of the nearest samples (equivalent to TSP), which leads to reduced performance. As $r$ increases, performance tends to degrade, eventually resembling that of random packing. For different models, while a smaller $r$ may benefit Llama2-7B by maintaining contextual relevance, Llama3-8B may require a larger $r$ to avoid the pitfalls of excessive similarity. This may occur because a more powerful model, being more sensitive to contextual reasoning, is more prone to performance degradation when exposed to overly similar or repetitive data.
6 Conclusion
We present TFP, a novel approach for packing samples during the SFT phase. This method exposes language models to relevant samples, allowing them to learn from relevant context and effectively adapt to few-shot evaluation. TFP is highly scalable and simple-to-implement, compatible with any existing SFT framework by simply altering the sequence of samples in the packing process. Our comprehensive evaluation demonstrates that TFP significantly enhances SFT performance for both text and tabular data, particularly in few-shot tasks. Additionally, TFP improves fairness in predictions. By adjusting the ratio within packs, TFP introduces samples with similar texts but varying sensitive attributes, which helps mitigate biases in LLMs without sacrificing accuracy.
Limitations
One limitation of this work is that we currently train on only one dataset at a time and evaluate using the corresponding evaluation methods for that dataset. To obtain usable LLMs, it is necessary to finetune them on a series of downstream tasks, which involves the selection and use of different datasets Liu et al. (2024); Ivison et al. (2023). Future research will explore the application of TFP in multi-task and multi-dataset settings, investigating the trade-offs of TFP in multi-task scenarios and potential issues in identifying relevant samples across multiple datasets.
Additionally, while TFP has made significant progress in fairness by modifying the ratio, many datasets lack annotations for sensitive attributes. The fairness improvements achieved by TFP largely depend on these annotations. Efficiently annotating datasets with sensitive attributes remains a challenge. A promising direction for future research could be exploring how to maintain balance within packs when such annotations are absent, further examining the relationship between TFP and responsible AI.
Moreover, we have not fully explored the relationship between providing related context within packs and other stages of training. Recent studies highlight the importance of maintaining internal knowledge consistency before and after SFT Ren et al. (2024); Yang et al. (2024). Earlier research on pretraining language models with related documents has shown promising results Staniszewski et al. (2024); Shi et al. (2024); Yasunaga et al. (2022); Yu et al. (2022). These methods either use metadata or retrieval techniques to group mutually relevant documents into long, coherent training examples. Our experiments indicate that providing context during SFT stages enhances in-context learning. We plan to explore integrating TFP with pretraining and in-context learning methods in future work.
Ethics Statement
Our study involves the development of a method for enhancing SFT, focusing on optimizing training efficiency and improving model performance. We also explored the potential of this method to improve fairness and mitigate bias. The dataset we used contains text that may be considered profane, vulgar, or offensive. The techniques and methodologies proposed in this paper are intended solely for research purposes and should not be applied to sensitive or high-risk domains without rigorous validation and oversight.
All the data used in this paper are publicly available and are used under the following licenses: MIT License, CC BY-NC 4.0 License, and CC0 1.0 License. All the LLMs used in this paper are used under the following licenses: Mistral AI Non-Production License, Llama 2 Community License, Meta Llama 3 Community License.
Appendix A Fairness Metrics
We introduce two commonly used definitions of group fairness metrics Wang et al. (2024a) for classification tasks. Suppose we have $n$ data samples $\{(X,Y,A)\}_{i=1}^{n}$ with features $X\in\mathcal{X}$, labels $Y\in\mathcal{Y}:=\{0,1\}$, and a sensitive attribute $A\in\{0,1\}$ drawn from the distribution $P_{XY}$. Note that the sensitive attribute $A$ is also included in the feature vector $X$. Let $f:\mathcal{X}\rightarrow\mathcal{Y}$ represent a machine learning model. We adopt the demographic parity difference metric $M_{dpd}$ to evaluate model prediction fairness:
$M_{dpd}=|P_{(X,Y,A)\sim P_{XY}}[f(X)=1\mid A=1]$ (1)
$-P_{(X,Y,A)\sim P_{XY}}[f(X)=1\mid A=0]|$
The demographic parity difference Zemel et al. (2013) measures the difference between the probability of positive predictions conditioned on the sensitive attribute $A=1$ and that conditioned on $A=0$. A large demographic parity difference $M_{dpd}$ indicates a significant prediction gap between the groups with $A=1$ and $A=0$, reflecting the model’s prediction unfairness. Since the demographic parity difference does not consider the ground truth label, we also use the metric of equalized odds difference $M_{eod}$ Hardt et al. (2016) to evaluate model prediction fairness:
$M_{eod}=\max\{M_{TP},M_{FP}\}$ (2)
where $M_{TP}$ denotes the true positive equalized odds difference:
$M_{TP}=|P_{(X,Y,A)\sim P_{XY}}[f(X)=1\mid Y=1,A=0]$ (3)
$-P_{(X,Y,A)\sim P_{XY}}[f(X)=1\mid Y=1,A=1]$ |||
and $M_{FP}$ denotes the false positive equalized odds difference:
$M_{FP}=|P_{(X,Y,A)\sim P_{XY}}[f(X)=1\mid Y=0,A=0]$ (4)
$-P_{(X,Y,A)\sim P_{XY}}[f(X)=1\mid Y=0,A=1]$ |||
A large equalized odds difference $M_{eod}$ indicates a significant prediction gap conditioned on different values of the sensitive attribute, indicating the model’s prediction unfairness.
Appendix B Experiment Details
We use the hyperparameters from the original framework for SFT, and the parameters of TFP are set in Section 5.
The prompts used for different datasets are as follows:
  • Adult dataset
    You are a helpful, pattern-following assistant. For each snippet of text, predict whether the person described has a good credit risk. The answer should be exactly ’yes’ or ’no’.
  • The Jigsaw Unintended Bias in Toxicity Classification task
    You are a helpful, pattern-following assistant. For each snippet of text, predict whether the text of the individual comment is harmless or toxic. The answer should be exactly ’harmless’ or ’toxic’.
  • Code task
    You are a professional programmer. Please provide the corresponding code based on the description of Human.
  • GSM8K and Alpaca
    Below is an instruction that describes a task. Write a response that appropriately completes the request.
Appendix C Evaluation Details
For the Adult dataset, we followed the setup from Wang et al. (2024a): We selected 200 samples to construct the test set. In both the test set and few-shot examples, the proportions of income and sex are balanced. We have converted the table data into a natural language format.
For the Jigsaw Unintended Bias in Toxicity Classification task, we select sensitive attributes similarly to the Adult dataset and use the same method to construct the test set.
The prompts used during evaluation are as follows:
  • The Jigsaw Unintended Bias in Toxicity Classification task
    You are a helpful, pattern-following assistant. For each snippet of text, predict whether the text of the individual comment is harmless or toxic. The answer should be exactly ’harmless’ or ’toxic’." + input_text + "Predict whether the above comment is harmless or toxic:
  • Adult dataset
    A person in 1996 has the following attributes: age 29, workclass Private, final weight 85572, education Bachelors, highest education level 13, marital status Married-civ-spouse, occupation Exec-managerial, relationship Wife, race White, sex Female, capital gain 0, capital loss 0, hours per week 42, native country United-States.predict whether the annual income of the person is greater than \$50k: